Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
berni - 22:59 Sunday 01 December 2019 (47828) Print this report
Operator Report - Afternoon shift

ITF found in Science mode.
It unlocked at 21:13:59 UTC and the problem seems to be due to missing data for all the ISYS signals, see plot1.
It relocked at second attempt, Science mode at 22:50 UTC.

Guard tours (times in UTC)

  • from 14:00 to 14:30;
  • from 16:10 to 16:40;
  • from 18:20 to 18:50;
  • from 20:10 to 20:40;

Air Conditioning
In the second plot you can see a trend of some relevant signals affected by the problem on the air conditioning machine occurred Friday night.

Images attached to this report
Comments to this report:
chiummo - 9:20 Monday 02 December 2019 (47830) Print this report

Follow-up of recovery after the HVAC failure:

Room temperature in the laser lab (and hosted benches) still experiencing the tail of a long transient. BPC horizontal corrections slowly drifting to compensate for the beam drifting on the sensors (not worrying, to be monitored). PMC seems to experience a little misalignment (Laser Amplifier power slightly increasing while PMC transmission is getting slightly worse), although from the reflected power the behavior is not vey clear. To monitor a possible increase of glitches related to the PSTAB if the misalignment of the PMC gets worse.

Overall input power for the interferometer decresed by 2%, to be re-checked after the steady state of the temperature is achieved.

For the moment, no urgent actions are needed from ISYS point of view, situation still under monitoring.

 

Images attached to this comment
Search Help
×

Warning

×